OpenStudy (anonymous):

A line has a slope of -3 and passes through the point (5, -1). What is the equation of the line?

OpenStudy (anonymous):

which form?

OpenStudy (anonymous):

you would use the y=mx=b formula and plus in the (-5,1) so you would write -1=-3(5)+b and solve for b. so for b you would get 14. so now your formula would be y=-3x+14!

OpenStudy (anonymous):

we have to use point slope form y-y1=m(x-x1) where our slope is m=-3 and our point is (5,-1) = (x1,y1) so x1=5 and y1=1 all you have to do is plug in x1 y1 and m into y-y1=m(x-x1)

OpenStudy (anonymous):

would it be positive 3...

OpenStudy (anonymous):

point slope form: y + 1 = -3(x-5)

OpenStudy (anonymous):

if the slope is -3 then the m would be -3. thats how i do it I'm not sure how your teacher wants